OpenStudy (anonymous):

\(\ \Huge (\sqrt{10})^3 ?\)

OpenStudy (anonymous):

\(\ \Huge ??? \)

hartnn (hartnn):

\(\sqrt{a}=a^{1/2}\)

hartnn (hartnn):

so \(\sqrt{10}^3=10^{3/2}=??\)

OpenStudy (anonymous):

You could cube 10 first then take the square root. So \[\sqrt{1000}\]

OpenStudy (anonymous):

Or, if they stay somewhat simple: \[ \left(\sqrt{10}\right)^3=\sqrt{10}\cdot\sqrt{10}\cdot\sqrt{10}=\left(\sqrt{10}\right)^2\sqrt{10}=10\sqrt{10} \]